Solve algebraically using only one variable. the length of a rectangle is two more than twice its width. If the area of the rectangle is 84, find the length and the width.

Questiond 2
Solve algebraically using only one variable. find three consecutive even intergers such that the product of the first integer and the third intger is equal to four less than ten times the second integer.

Answer by vertciel(183)

1.
Express in terms of variables:
Width = w
Length = 2 + 2w
LW = Area of rectangle
(2 + 2w)w = 84
Find w and substitute.

2. Denote the consecutive even integers as n, n + 2, and n + 4.
n(n + 4) = 10(n + 2)
n^2 - 6n - 16 = 0
Solve n, using factoring.